Find a vector normal to the surface z = el-y at the point (1,0, e).
Answer #1

I am uploading hand written solution for this question. I hope it will help you to understand. Please like and comment if you found helpful.

Given - Z= e2-4² f(x, y, z) = el-y-z =0 the gradient of fa, y, z) at point x, y, z ý a vector normal to the surface at this p23 24 25 26 21 20 L of (1,0,0) = (de, 0, -1) | Hence the vector honal to the surface z=ety at poin (1,0, e į = (20, 0, -1) Au